OpenStudy (anonymous):

Factor completely: 8x4 + 4x3 - 24x2 4x2(2x2 + x – 6) 4x2(2x – 3)(x + 2) 4(2x4 + x3 - 6x2) prime

OpenStudy (amistre64):

2(6) = 12; +4-3 = 1 4/2; 3/2 x+2 2x-3 is what i get to

OpenStudy (anonymous):

la, if u could just help me finish this times test and then go back and help me actually understand this stuff thatd be greatt!!! (:

OpenStudy (anonymous):

4x2(2x – 3)(x + 2)

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

the final answer is \[4x ^{2}(2x-3)(x+2)\]

OpenStudy (anonymous):

okay the next problem is Factor completely: 4x4 - 64 4(x4 - 16) 4(x2 + 4)(x2 - 4) 4(x2 + 4)(x + 2)(x – 2) prime

jimthompson5910 (jim_thompson5910):

\[4x^4-64=4 (x^2+4)(x+2)(x-2) \] by use of difference of squares

OpenStudy (anonymous):

4(x2 + 4)(x + 2)(x – 2)

OpenStudy (anonymous):

okat two more (:

OpenStudy (anonymous):

One of the solutions to the quadratic equation x2 - 13x + 30 is 3. What is the other solution?

jimthompson5910 (jim_thompson5910):

10

OpenStudy (anonymous):

10

jimthompson5910 (jim_thompson5910):

x^2-13x+30=(x-10)(x-3)

OpenStudy (anonymous):

One of the solutions to the quadratic equation 2x2 - 9x – 5 is -1/2. What is the other solution?

OpenStudy (anonymous):

im gonna go back and have you guys explain this to me after i submit this test

jimthompson5910 (jim_thompson5910):

other solution is 5

OpenStudy (anonymous):

Solve by using the quadratic formula. x2 + 4x + 1 = 0.

OpenStudy (anonymous):

-1/2

jimthompson5910 (jim_thompson5910):

solutions are \[x=-2+\sqrt{3}\] or \[x=-2-\sqrt{3}\]

OpenStudy (anonymous):

Solve by using the quadratic formula. 3x2 - 4x = 2.

jimthompson5910 (jim_thompson5910):

Solutions are \[x=\frac{2+\sqrt{10}}{3}\] or \[x=\frac{2-\sqrt{10}}{3}\]